# Combinatorial Auctions without a Numeraire: The Case of Blockchain Trade-Intent Auctions

### v1

@@ Abstract
Blockchain trade intent auctions currently intermediate approximately USD 5 billion monthly. Due to production complementarities, the auction is combinatorial: when multiple trade intents from different traders are auctioned off simultaneously, a bidder (here called solver) can generate additional efficiencies by winning a batch of multiple trade intents. However, unlike other combinatorial auctions studied in the literature, the auction has no numeraire. Fairness is a concern as the efficiencies from batching cannot be easily shared between traders. We formalize this problem and study the most commonly used auction formats: batch auctions and multiple simultaneous auctions. We also propose a novel fair combinatorial auction that combines batch auction and multiple simultaneous auctions: solvers submit individual-trade bids and batched bids, but batched bids are considered only if they are better for all traders relative to the outcome of multiple simultaneous auctions (constructed using the individual-trade bids). We find a trade-off between the fairness guarantees provided by the auction (i.e., the minimum each trader can expect to receive) and the expected value of the assets returned to the traders. Also, the amount that each trader receives in the equilibrium of the fair combinatorial auction may be higher or lower than what they receive in the equilibrium of the simultaneous auctions used as a benchmark for fairness.
August 22, 2024  (first draft, all comments are welcomed)
@@ 1 Introduction
Permissionless blockchains like Ethereum allow anyone to create blockchain-based applications at minimal or no cost, including blockchain-based financial markets. As a result, there are now dozens of such markets, each with a different mechanism, fee structure, and liquidity. The resulting market fragmentation generates complexity for users because executing a trade most efficiently may require accessing multiple marketplaces, often via intermediate assets. Also, these complex executions expose users to “sandwich attack”, in which an attacker front-runs a victim’s trade with a trade in the same direction and then back-runs it with a trade in the opposite direction. Doing so allows the attacker to “buy cheap” and “sell expensive” while forcing the victim to trade at less favorable terms.
@@
Trade-intent auctions emerged as a response to this complexity and these risks. A trade intent is a “partial blockchain transaction” that specifies an amount of a sell asset and a buy asset (or an amount of a buy asset and a sell asset) without specifying how the transaction should be executed. The right to complete this partial transaction is auctioned off to specialized entities called solvers, who compete to provide the best possible execution to each trade intent. The first notable trade-intent auction was CoW Swap’s batch auction, in which a group of trade intents are actioned off simultaneously, solvers submit a single bid that may contain multiple trade intents, and the auctioneer selects a single winner. Later entrants, such as 1inch (1inch fusion) and Uniswap (UniswapX), instead hold Dutch or descending price auctions, one per each trade intent. Overall, trade-intent auctions currently intermediate approximately USD 5 billion monthly. [^1] More precisely, the total volume of trade-intent auctions on Ethereum was USD 5.9 billion in April 2024, USD 7.7 billion in March 2024, USD 4.9 billion in May 2024, USD 4.2 billion in June 2024, USD 5.2 billion in July 2024 (source: https://dune.com/queries/3058985/5091785). For comparison, the widely studied UK 3G spectrum auction held in 2001 generated USD 34 billion (or approx USD 60 billion at current prices).
@@
From the solvers’ perspective, competing in the auction requires a balancing act between specialization in certain trade intents and the ability to exploit complementarities between trade intents. To specialize, some solvers may develop sophisticated algorithms to access public marketplaces, while others may work closely with private market makers. To exploit complementarities between trade intents, solvers may find coincidences of wants allowing traders to trade with each other, sometimes in complex ring trades or intermediate steps of multi-step trades. Additionally, because blockchain fees (i.e., gas costs) are largely independent of trade volume, two orders may save on fees if executed together. Solvers also face the risk of sandwich attacks and employ complex strategies to mitigate them.
@@
From the theoretical viewpoint, trade-intent auctions are combinatorial because how a solver can best execute a trade depends on the other trades the solver also executes. Similarly to most combinatorial auctions, the goal of the auction designer is to balance the benefit of specialization (i.e., allocating each trade intent to the solver who is best at that trade intent) with the additional efficiencies from batching (i.e., allocating multiple orders to the same solver). But unlike most combinatorial auctions, the assets exchanged in the auctions are illiquid: the premise of the auction is that even if each asset has a notional market value, exchanging them is subject to frictions and fees. Hence, combinatorial auctions lack a common numeraire to reallocate the efficiencies from batching between traders.
@@
We then use the model to compare the most commonly used auction mechanisms: batch action and simultaneous standard auctions (as in Gentry et al., 2019). At a superficial level, one could think that simultaneous standard auctions are better at exploiting the benefit of specialization because different solvers may win different orders. Batch auctions, instead, are better at exploiting the complementarities between various orders. Hence, simultaneous standard auctions should be preferred when the benefit of specialization is large, while batch auctions should be preferred when the complementarities between different orders are strong. This reasoning, however, misses an important element: when the benefit of specialization is strong, competition is low in the simultaneous standard auctions. In other words, specialization implies that, in each standard auction, there is a “leader” who can win the auction easily — that is, can win while returning little to the traders. On the other hand, batching forces all solvers to compete with each other: each “leader” in a given type of trade needs to outbid the other leaders to win any trade. Hence, solvers have a higher incentive to bid (and return a higher fraction of the assets produced) in the batch auction relative to the simultaneous standard auctions. At the same time, the batch auction has multiple equilibria because the solvers’ payoff only depends on the value of the assets returned to the traders, not how this value is shared. Therefore, the benefit of the batch auction may accrue disproportionally to one trader, leaving the other worse off relative to the simultaneous standard auctions. At an intuitive level, such an outcome seems unfair. [^2] CoW Swap implementation of the batch auction tackles this problem by adding a period to the game: after the assets are returned to the users, CoW Swap checks the solver’s execution against the prices on the most popular blockchain-based financial marketplaces. If the solver’s execution of an order falls below such reference, CoW Swap asks the solver to reimburse the user. The main drawback is that the benchmark to define whether the execution of an order is fair is determined by CoW Swap and does not incorporate solvers’ knowledge.
@@
Motivated by this observation, we then introduce a definition of fairness inspired by the literature on cooperative game theory. A weak notion of fairness is that traders are batched together if and only if they all benefit relative to a reference outcome. In cooperative game theory, this reference outcome is the players’ “outside options”: the highest possible payoff each player can achieve without the collaboration of the other players. In the case of the combinatorial auction, instead, we assume that the reference outcome is the auction’s outcome when the batched bids are removed. [^3] If multiple batched bids are better than the reference outcome for both players, then we can strengthen the notion of fairness, for example, by ranking different allocations by the product of the extra utility relative to the benchmark. As it will become clear later, our main results continue to hold for any strengthening of the notion of fairness. The reference outcome is, therefore, constructed as part of the mechanism and depends on the solvers’ bids. For ease of comparison with the previously discussed mechanisms, when solving for the equilibrium of the fair combinatorial auction, we assume that the auction can be decomposed into two individual trade auctions, which can be in first price or second price, and a batched auction in which the winner “pays its bid” (i.e., it returns to the traders the amounts specified in its batched bid). Fairness implies that the auctioneer ignores a batched bid whenever it delivers less than the outcome of the two individual trade auctions, and it always chooses a batched bid whenever there is one that is fair.
@@
If the individual-trade auctions are in second price, then each solver has limited ability to influence the fairness benchmark: a solver who bids higher than the opponent in the individual orders can never increase the fairness benchmark, and a solver who bids lower than the opponent can increase the fairness benchmark just up to the opponent’s bids. As a consequence, in equilibrium, the auction may not provide any fairness guarantee. For example, if both solvers set their individual-trade bids to zero (which is equivalent to not bidding on individual trade), then the equilibrium is always batching, and the only additional “fairness” guarantee provided by auction is that each trader should receive a strictly positive amount of the asset demanded. In practice, the two effects described earlier are mute, and the outcome of this fair combinatorial auction is identical to that of the batched auction.
If, instead, the individual-trade auctions are in first price, then each solver can manipulate the fairness benchmark with his individual-trade bids. The two effects described earlier are present, and solvers will bid higher in the individual-trade auction but lower in the batched bid. In equilibrium, we find three possible regimes:
  • Specialization: if, on each trade, the difference in solvers’ productivity is very large, then each solver will disqualify the opponent’s batched bid as unfair. The outcome is specialization, and each trader receives the assets that the worse solver at his order would provide when matched with both orders. This last part is important because it implies that traders receive more than what they would receive in a simultaneous standard auction (as those held by 1 Inch and Uniswap).
  • Competitive Batching: when the benefit of batching is large relative to the traders’ productivity, then the equilibrium is similar to that of the batch auction, with one important difference: the equilibrium is unique, and each trader receives at least the assets that the worst solver at his order would provide when matched with both orders.
@@
Hence, relative to the fair combinatorial auction in which the individual-trade auctions are in the second price, when the individual-trade auctions are in first price two things happen. First, the auction provides stronger fairness guarantees. At the same time, we show that the market value of the assets delivered to the users is lower, strictly so if the outcome is specialization or uncompetitive batching.
The comparison between the two mechanisms, therefore, illustrates two important points. First, in equilibrium, higher fairness guarantees come at the expense of the market value of the assets delivered. Second, the minimum that each trader receives in equilibrium may be very different from the equilibrium of the mechanism used as a benchmark for fairness. For example, when the mechanism used as a reference for fairness is two second-price auctions, one of the traders may earn less in the equilibrium of the fair combinatorial auction than in the equilibrium of the second-price auctions. If the reference for fairness is two first-price auctions, then both traders always earn more in the equilibrium of the fair combinatorial auction than in the equilibrium of the first-price auctions.
@@ 1.1 Relevant literature
This paper is related to several strands of literature. Che and Gale (1998) study a single auction with financially constrained bidders and show that a first-price auction generates higher revenues than a second-price auction. The intuition is that bids are lower and the constraint is less likely to bind in a first-price auction than in a second-price auction. Their work is related to our paper because illiquidity also imposes a constraint on feasible bids. The difference is that, in the case of illiquidity, such a constraint depends on the assignment of trade intents to solvers and hence is relevant only in a combinatorial problem. [^4] Dobzinski et al. (2012) study multiple simultaneous auctions with financially constrained bidders. In their paper, however, there is a single budget constraint over the sum of the possible bids, which is quite different from the feasibility constraint generated by illiquidity. We are also related to the literature studying multiple simultaneous auctions, such as, for example, Milgrom (2000), Krishna and Rosenthal (1996), Goeree and Lien (2014), and Gentry et al. (2019). In particular, Gentry et al. (2019) studies simultaneous sealed bit auctions with complementarities, corresponding to one of the benchmarks we consider (modulo the feasibility constraint we introduce).
@@
Finally, our paper relates to the literature studying how frictions in the ability to reallocate utility between players affect the equilibrium of cooperative games. In this literature, games in which it is possible to frictionlessly reallocate utility between players are called TU (Transferable Utility) games, and those in which this frictionless reallocation is not possible are called NTU (Non-TU) games. For example, it is well known that several cooperative game theory concepts are equivalent under TU but different under NTU. Also, the equilibrium in matching games with TU may change under NTU (Becker, 1973, Legros and Newman, 2007). The main difference is that, in cooperative games, the reference for fairness (i.e., the outside options) is common knowledge. In contrast, here, this reference needs to be elicited as part of the mechanism.
inlineinlinetodo: inline “While interest in the issues represented by the non-transferable case is both longstanding and lively (see for instance Farrell-Scotchmer, 1988 on production in partnerships; Rosenzweig-Stark, 1989 on risk sharing in households; and more recently, Lazear, 2000 on incentive schemes for workers; Ackerberg-Botticini, 2002 on sharecropping; and Chiappori-Salanié, 2003 on the empirics of contracts)”,
@@ 2 The model
  • If solver 1 is matched with both orders, it produces $g\cdot\beta$ units of asset B and $g\cdot\underline{\delta}$ units of asset D for $g\geq 1$. If solver 2 is matched with both orders, it produces $g\cdot\underline{\beta}$ units of asset B and $g\cdot\delta$ units of asset D. We assume $g\geq 1$ to capture the additional efficiencies when a solver executes both orders.
Hence, each solver is specialized in a given order: solver 1 is better at order 1, and solver 2 is better at order 2. Note also that there is a single way to execute both orders: when matched with both orders, a solver cannot choose to generate more of a given asset at the expense of the other asset (of course, as we’ll see later, when bidding a solver can return more of a given asset at the expense of the other asset). In Section 6.2, we relax this assumption by introducing the choice of how to execute both orders. Finally, the parameters, $\underline{\beta}$, $\overline{\beta}$, $\underline{\delta}$, $\overline{\delta}$, and $g$ are common knowledge among the players but are unknown to the designer, who therefore cannot condition the mechanism on their values (see below).
@@
where the first entry corresponds to the solver matched with the first order and similarly for the second entry. Call $x_{1}(B)$ and $x_{1}(D)$ the amount of token $B$ and $D$ returned by solver 1, and similarly define $x_{2}(B)$ and $x_{2}(D)$. Our main assumption is that solvers cannot return to the traders more than the sum of the assets than what they produce and their inventory: [^5] Note the similarity between the feasibility constraint and a budget constraint (already mentioned in the literature review when discussing Che and Gale, 1998): they both impose a constraint on the bids that can be paid. However, in Assumption 1, such constraint is a function of $\mu$; that is, it is contingent on the trades the solver wins.
@@
We can now write the solvers’ payoff as a function of $\mu$, $x_{1}(B)$, $x_{1}(D)$, $x_{2}(B)$, and $x_{2}(D)$ such that the feasibility constraint hold.

### v2

@@ Abstract
Blockchain trade intent auctions currently intermediate approximately USD 5 billion monthly. Due to production complementarities, the auction is combinatorial: when multiple trade intents from different traders are auctioned off simultaneously, a bidder (here called solver) can generate additional efficiencies by winning a batch of multiple trade intents. However, sharing these additional efficiencies between traders is problematic: because of market frictions and fees (solvers’ private information), the auctioneer does not know how much each trader would have received had its trade been auctioned off individually. We formalize this problem and study the most commonly used auction formats: batch auctions and multiple simultaneous auctions. We also propose a novel fair combinatorial auction that combines batch auction and multiple simultaneous auctions: solvers submit individual-trade bids and batched bids, but batched bids are considered only if they are better for all traders relative to the outcome of the simultaneous auctions constructed using the individual-trade bids. We find a trade-off between the fairness guarantees provided in equilibrium by the auction (i.e., the minimum each trader can expect to receive) and the expected value of the assets returned to the traders. Also, the amount that each trader receives in the equilibrium of the fair combinatorial auction may be higher or lower than what they receive in the equilibrium of the simultaneous auctions used as a benchmark for fairness.
This version: December 2, 2024
@@ 1 Introduction
Permissionless blockchains such as Ethereum allow anyone to create blockchain-based applications at minimal or no cost, including blockchain-based financial markets. As a result, there are now dozens of such markets, each with a different mechanism, fee structure, and liquidity. The resulting market fragmentation generates complexity for users because executing a trade most efficiently may require accessing multiple marketplaces, often via intermediate assets. Also, these complex executions expose users to “sandwich attack”, in which an attacker front-runs a victim’s trade with a trade in the same direction and then back-runs it with a trade in the opposite direction. Doing so allows the attacker to “buy cheap” and “sell expensive” while forcing the victim to trade at less favorable terms.
@@
Trade-intent auctions emerged as a response to this complexity and these risks. A trade intent is a “partial blockchain transaction” that specifies an amount of a sell asset and a buy asset (or an amount of a buy asset and a sell asset) without specifying how the transaction should be executed. The right to complete this partial transaction is auctioned off to specialized entities called solvers, who bid by proposing the best possible execution to each trade intent. The oldest trade-intent auction is the batch auction held by CoW Protocol (whose main user interface is CoW Swap), in which a group of trade intents are auctioned off simultaneously, solvers submit a single bid that may contain multiple trade intents, and the auctioneer selects a single winner. Later entrants, such as 1inch (1inch fusion) and Uniswap (UniswapX), instead hold a descending price (or Dutch) auction for each trade intent. Trade-intent auctions are repeated every few seconds and involve between 10 and 25 solvers. At the time of writing, they intermediate more than USD 5 billion monthly. [^1] Source: https://dune.com/queries/3058985/5091785. For comparison, the widely studied UK 3G spectrum auction held in 2001 generated USD 34 billion (or approx USD 60 billion at current prices).
An important characteristic of trade-intent auctions is that there is no numeraire: users (who are anonymous and may live anywhere in the world) can demand to exchange any asset for any other asset. This characteristic makes them similar to foreign exchange trading desks but for one crucial aspect: the number of assets traded in trade intent auctions is orders of magnitude larger than that in foreign exchange markets, and this difference is further magnified if we consider asset pairs. For example, in October 2024, CoW Protocol facilitated trades between 2.500 different assets on Ethereum, with the top 5 asset pairs accounting for 28% of the total monthly volume. At the same time, there are approximately 100 different currencies traded in Foreign Exchange markets, with the top 5 pairs accounting for about 65% of total volume (and all these pairs involve the US Dollar).
@@
For solvers, competing in a trade-intent auction requires a balancing act between specialization in certain trade intents and the ability to exploit complementarities between trade intents. To specialize, some solvers may develop sophisticated algorithms to access public marketplaces, others may work closely with private market makers, while others may specialize in assets requiring special operations. [^2] Remember that blockchain-based assets such as cryptocurrencies and crypto-tokens are “programmable money,” and this programming can be quite arbitrary. For example, there are assets that can be redeemed for some other underlying assets (which can then be traded) or that lose value every time they are exchanged. To exploit complementarities between trade intents, solvers may find coincidences of wants allowing traders to trade with each other, sometimes in complex ring trades or intermediate steps of multi-step trades. Additionally, because blockchain fees (i.e., gas costs) are largely independent of trade volume, two orders may save on fees if executed together. Solvers also face the risk of sandwich attacks and employ complex strategies to mitigate them.
@@
From the theoretical viewpoint, trade-intent auctions are combinatorial because how a solver can best execute a trade depends on the other trades the solver also executes. Similarly to most combinatorial auctions, the goal of the auction designer is to balance the benefit of specialization (i.e., allocating each trade intent to the solver who is best at that trade intent) with the additional efficiencies from batching (i.e., allocating multiple orders to the same solver). But unlike most combinatorial auctions, the assets exchanged in the auctions are illiquid: the premise of the auction is that even if each asset has a notional market value, exchanging them is subject to frictions and fees. Together with the fact that there is no common numeraire, this implies that trade-intent auctions lack a way to efficiently share the benefit of batching between the traders.
@@
We then use the model to compare the most commonly used auction mechanisms: batch action and simultaneous standard auctions (as in Gentry et al., 2019). At a superficial level, one could think that simultaneous standard auctions are better at exploiting the benefit of specialization because different solvers may win different orders. Batch auctions, instead, are better at exploiting the complementarities between various orders. Hence, simultaneous standard auctions should be preferred when the benefit of specialization is large, while batch auctions should be preferred when the complementarities between different orders are strong. This reasoning, however, misses an important element: when the benefit of specialization is strong, competition is low in the simultaneous standard auctions. In other words, specialization implies that, in each standard auction, there is a “leader” who can win the auction easily — that is, can win while returning little to the traders. On the other hand, batching forces all solvers to compete with each other: each “leader” in a given type of trade needs to outbid the other leaders to win any trade. Hence, solvers have a higher incentive to bid (and return a higher fraction of the assets produced) in the batch auction relative to the simultaneous standard auctions. At the same time, the batch auction has multiple equilibria because the solvers’ payoff only depends on the value of the assets returned to the traders, not how this value is shared. Therefore, the benefit of the batch auction may accrue disproportionally to one trader, leaving the other worse off relative to the simultaneous standard auctions. At an intuitive level, such an outcome is unfair. [^3] CoW Protocol implementation of the batch auction tackles this problem by adding a period to the game: after the assets are returned to the users, CoW Protocol checks the solver’s execution against a counterfactual execution that uses exclusively the most popular blockchain-based financial marketplaces. If the solver’s execution of an order falls below such reference, CoW Protocol asks the solver to reimburse the user. The main drawback is that the benchmark to define whether the execution of an order is fair is determined by CoW Protocol and does not incorporate solvers’ knowledge.
@@
Motivated by this observation, we then introduce the fair combinatorial auction, in which solvers can bid on individual trades and also on batches of trades. Inspired by the literature on cooperative game theory, the auction is fair because traders are batched together if and only if they all benefit relative to a reference outcome. In cooperative game theory, this reference outcome is the players’ “outside options”: the highest possible payoff each player can achieve without the collaboration of the other players. In the case of the combinatorial auction, instead, we assume that the reference outcome is the auction’s outcome when only bids on individual orders are considered. [^4] If multiple batched bids are better than the reference outcome for both players, then we can strengthen the notion of fairness, for example, by ranking different allocations by the product of the extra utility relative to the benchmark. As it will become clear later, our main results continue to hold for any strengthening of the notion of fairness. The reference outcome is, therefore, constructed as part of the mechanism and depends on the solvers’ bids. For ease of comparison with the previously discussed mechanisms, when solving for the equilibrium of the fair combinatorial auction, we assume that the auction can be decomposed into two individual trade auctions, which can be in first price or second price, and a batched auction in which the winner “pays its bid” (i.e., it returns to the traders the amounts specified in its batched bid). Fairness implies that the auctioneer ignores a batched bid whenever it delivers less than the outcome of the two individual trade auctions, and it always chooses a batched bid whenever there is one that is fair.
@@
If the individual-trade auctions are in second price, then each solver has limited ability to influence the fairness benchmark: a solver who bids higher than the opponent in the individual orders can never increase the fairness benchmark, and a solver who bids lower than the opponent can increase the fairness benchmark just up to the opponent’s bids. As a consequence, in equilibrium, the auction may not provide any fairness guarantee. For example, if both solvers set their individual-trade bids to zero (which is equivalent to not bidding on individual trade), then the equilibrium is always batching and the auction does not provide any meaningful fairness guarantees. The two effects described earlier are mute, and the set of outcomes of this fair combinatorial auction is identical to that of the batch auction.
If, instead, the individual-trade auctions are in first price, then each solver can manipulate the fairness benchmark with his bids. The two effects described earlier are present, and solvers will bid higher in the individual-trade auction but lower in the batched bid. In equilibrium, we find three possible regimes:
  • Specialization: if, on each trade, the difference in solvers’ productivity is very large, then each solver will disqualify the opponent’s batched bid as unfair. The outcome is specialization, and each trader receives the assets that the worse solver at his order would provide when matched with both orders. This last part is important because, even if the outcome is specialization, traders receive more than what they would receive in a simultaneous standard auction.
  • Competitive Batching: when the benefit of batching is large relative to the traders’ productivity, then the equilibrium is similar to that of the batch auction, with one important difference: each trader receives at least the assets that the worst solver at his order would provide when matched with both orders.
@@
Hence, the auction provides strong fairness guarantees, as both traders receive more than what they would receive in the simultaneous standard auctions. At the same time, we show that the market value of the assets delivered to the users is lower than in the batch auction, strictly so if the outcome is specialization or uncompetitive batching.
The comparison between the two cases (i.e., the fair combinatorial auction in which the individual-trade auctions are in first price vs. in second price) illustrates two important points. First, in equilibrium, stronger fairness guarantees come at the expense of the market value of the assets delivered. Second, the minimum that each trader receives in equilibrium may be very different from the equilibrium of the mechanism used as a benchmark for fairness. For example, when the mechanism used as a reference for fairness is two second-price auctions, one of the traders may earn less in the equilibrium of the fair combinatorial auction than in the equilibrium of the second-price auctions. If the reference for fairness is two first-price auctions, then both traders always earn more in the equilibrium of the fair combinatorial auction than in the equilibrium of the first-price auctions.
@@ 1.1 Relevant literature
The problem we study is related to several strands of literature. It is similar to multi-units or combinatorial assignment problems because there is no numeraire (see Budish, 2011, Budish and Cantillon, 2012). However, assignment problems exclude the possibility of monetary transfers. In trade intent auctions, instead, there are many types of monetary transfers, and some users are willing to accept some transfers but not others. Note also that the literature studying assignment problems has proposed several definitions of fairness, aiming at capturing whether, in equilibrium, an agent is “envious” of a different agent’s allocation. Fairness is, therefore, defined by comparing agents within the same mechanism. Instead, we introduce a novel notion of fairness that compares the outcome of a mechanism with that of a different mechanism.
Our definition of fairness is heavily inspired by the literature on cooperative game theory. If we think of the players’ outside options as holding separate trade-by-trade auctions and “cooperation” as batching the two trades together, then cooperation can emerge only if each player earns as much by cooperating as he/she would earn in his/her outside option. That literature has also studied extensively the differences between games in which the frictionless reallocation of utility between players is possible (so-called Transferable Utility or TU games) and those in which it is not (so-called Non-TU or NTU games). In particular, if solvers’ production function was common knowledge, then the problem we study would be a matching problem with non-transferrable utility, similar to that studied in Legros and Newman, 2007 who allow for the reallocation of utility via transfers, but with frictions. The main difference is that, in cooperative games, the production function and the reference for fairness (i.e., the outside options) are common knowledge. In contrast, here, both the solvers’ technology and the reference for fairness need to be elicited as part of the mechanism.
We are, therefore, related to the literature on auctions and mechanism design. In particular, Che and Gale (1998) study a single auction with financially constrained bidders and show that a first-price auction generates higher revenues than a second-price auction. The intuition is that bids are lower and the constraint is less likely to bind in a first-price auction than in a second-price auction. Their work is related to our paper because illiquidity also imposes a constraint on feasible bids. The difference is that, in the case of illiquidity, such a constraint depends on the assignment of trade intents to solvers and hence is relevant only in a combinatorial problem. [^5] Dobzinski et al. (2012) study multiple simultaneous auctions with financially constrained bidders. In their paper, however, there is a single budget constraint over the sum of the possible bids, which is quite different from the feasibility constraint generated by illiquidity. We are also related to the literature studying multiple simultaneous auctions, such as, for example, Milgrom (2000), Krishna and Rosenthal (1996), Goeree and Lien (2014), and Gentry et al. (2019). In particular, Gentry et al. (2019) studies simultaneous sealed bit auctions with complementarities, corresponding to one of the benchmarks we consider (modulo the feasibility constraint we introduce).
@@ 2 The model
  • If solver 1 is matched with both orders, it produces $g\cdot\beta$ units of asset B and $g\cdot\underline{\delta}$ units of asset D. If solver 2 is matched with both orders, it produces $g\cdot\underline{\beta}$ units of asset B and $g\cdot\delta$ units of asset D. We assume $g\geq 1$ to capture the additional efficiencies when a solver executes both orders.
Hence, each solver is specialized in a given order: solver 1 is better at order 1, and solver 2 is better at order 2. Note also that we abstract away from the choice of how to execute the orders, that is, when matched with both orders, a solver cannot choose to generate more of a given asset at the expense of the other asset (of course, as we’ll see later, when bidding a solver can return more of a given asset at the expense of the other asset). In Section 6.3, we relax this assumption by introducing the choice of how to execute both orders. Finally, the parameters, $\underline{\beta}$, $\overline{\beta}$, $\underline{\delta}$, $\overline{\delta}$, and $g$ are common knowledge among the players but are unknown to the designer, who therefore cannot condition the mechanism on their values (see below). Instead, $\beta$ is solver 1’s private information and $\delta$ is solver 2’s private information.
@@
where the first entry corresponds to the solver matched with the first order and the second entry corresponds to the solver matched with the second order. Call $x_{1}(B)$ and $x_{1}(D)$ the amount of token $B$ and $D$ returned by solver 1, and similarly define $x_{2}(B)$ and $x_{2}(D)$. Our main assumption is that solvers cannot return to the traders more than the sum of what they can produce and their inventory: [^6] Note the similarity between the feasibility constraint and a budget constraint (already mentioned in the literature review when discussing Che and Gale, 1998): they both impose a constraint on the bids that can be paid. However, in Assumption 1, such constraint is a function of $\mu$; that is, it is contingent on the trades the solver wins.
@@
We can now write the solvers’ payoff as a function of $\mu$, $x_{1}(B)$, $x_{1}(D)$, $x_{2}(B)$, and $x_{2}(D)$.